Two small positively charged spheres have a combined charge of 5.0x10-5 C. If each sphere is repelled from the other by an electrostatic force of 1.0 N when the sphere are 2.0 m apart, what is the charge on each sphere?


1
Expert's answer
2021-10-25T17:54:25-0400
"F=k\frac{q(Q-q)}{r^2}\\1=9\cdot10^{9}\frac{q(5\cdot10^{-5}-q)}{2^2}\\q=3.8\cdot10^{-5}C\\Q-q=1.6\cdot10^{-5}C"